OpenStudy (anonymous):

Chris’s hike started at an elevation of –12 m. His hike finished at an elevation of 235 m. How much did Chris’s elevation change?

OpenStudy (igreen):

Just add.. 235 + -12 = 223. So Chris's elevation changed by 223m.

OpenStudy (texaschic101):

I am not sure about that....when you add 12 to -12, it equals 0, then there is 235...so wouldn't it be 235 + 12 = 247m

OpenStudy (texaschic101):

@iGreen

OpenStudy (igreen):

Well if he starts at -12..it takes 12m just to get to zero and on top of that you have 235..so yeah I think you're right..

OpenStudy (texaschic101):

I wasn't sure there for a second...but confirmation is always good

OpenStudy (perl):

235 - (-12) = 235 + 12